# Boosting Data Analytics with Synthetic Volume Expansion

### 07 Jan 2024 v2

@@ Abstract
Synthetic data generation, a cornerstone of Generative Artificial Intelligence (GAI), signifies a paradigm shift in data science by addressing data scarcity and privacy while enabling unprecedented performance. As synthetic data gains prominence, questions arise concerning the accuracy of statistical methods when applied to synthetic data compared to raw data. This article introduces the Synthetic Data Generation for Analytics (Syn) framework. This framework employs statistical methods on high-fidelity synthetic data generated by advanced models such as tabular diffusion and Generative Pre-trained Transformer (GPT) models. These models, trained on raw data, are further enhanced with insights from pertinent studies through knowledge transfer. A significant discovery within this framework is the generational effect: the error of a statistical method on synthetic data initially diminishes with additional synthetic data but may eventually increase or plateau. This phenomenon, rooted in the complexities of replicating raw data distributions, highlights a “reflection point” — an optimal threshold in the size of synthetic data determined by specific error metrics. Through three case studies — sentiment analysis of texts, predictive modeling of structured data, and inference in tabular data — we demonstrate the effectiveness of this framework over traditional ones. We underline its potential to amplify various statistical methods, including gradient boosting for prediction and hypothesis testing, thereby underscoring the transformative potential of synthetic data generation in data science.
@@ 1 Introduction
The advent of synthetic data generation, fueled by generative artificial intelligence (AI), has shifted data analytics towards a more synthetic data-centric approach. According to Gartner, 60% of the data utilized in AI and analytics projects will be synthetically generated by 2024, and synthetic data will surpass real data in AI models by 2030 (Gartner, 2022; Eastwood, 2023). This paradigm shift challenges the traditional practice, which exclusively relies on raw data. Synthetic data, mirroring real-world scenarios, presents a viable alternative to the challenges posed by data collection, sharing, and analysis within limited data environments.
Synthetic data confers two primary advantages on data analytics (Walonoski et al., 2018; Topol, 2019). First, it alleviates data scarcity and addresses privacy concerns (Ghalebikesabi et al., 2023). When crafted to emulate the raw data distribution (Bi and Shen, 2023; Shen, Bi and Shen, 2022; Xie et al., 2018), sharing synthetic data comes with minimal risk of exposing sensitive raw data. The significance of such data is accentuated in downstream or subsequent analyses, as exemplified by a COVID-19 study (El Emam et al., 2021). Second, synthetic data enables training in real-world scenarios like autonomous driving (Sun et al., 2020; Liu, Shen and Shen, 2023), negating the necessity for expensive experiments. Moreover, it proves invaluable for approximating the distributions of test statistics using Monte Carlo methods through repeated numerical experiments (Liu, Shen and Shen, 2023).
@@
This paper introduces the Synthetic Data Generation for Analytics (Syn) framework, designed to bolster the precision of any statistical methods using high-fidelity synthetic data that closely mirrors raw data, thereby harnessing the anticipated advantages. Generative models produce synthetic data by training on raw data and enrichment with insights from related studies via knowledge transfer. The Syn framework employs an array of generative models suited for various domains: image diffusion (Sohl-Dickstein et al., 2015; Wei et al., 2021), text diffusion models (Ho, Jain and Abbeel, 2020), text-to-image diffusion models (Zhang et al., 2023), time-series diffusion models (Lin et al., 2023), spatio-temporal diffusion models (Yuan et al., 2023), and tabular diffusion models (Kotelnikov et al., 2023; Zheng and Charoenphakdee, 2022; Kim, Lee and Park, 2022; Lee, Kim and Park, 2023). Moreover, Syn embraces advanced models such as the Reversible Generative Models (Kingma and Dhariwal, 2018). These flow-based models capture the raw data distribution and can estimate both conditional and marginal distributions. Central to our Syn exploration is this pivotal issue: Can high-fidelity synthetic data enhance the efficacy of statistical methods solely reliant on raw data? If so, how may we implement such enhancements? Recent research offers diverging viewpoints on this issue. In some cases, synthetic X-ray images improve the accuracy of machine learning models (Gao et al., 2023), whereas, in others, training on synthetic data may compromise performance for some machine learning models (Kotelnikov et al., 2023).
@@
Synthetic data holds immense potential for enhancing data analytics. When generated accurately, this high-fidelity data can boost the accuracy of a statistical method by expanding the sample size of raw data. However, a significant caveat exists: low-fidelity synthetic data could yield unreliable outcomes. Often, a generational effect emerges, whereby as the size of synthetic data grows, the precision gain might diminish or even plateau. This phenomenon has been exemplified in our case study focusing on structured data prediction as discussed in Section 4.2. This challenge arises from generation errors or discrepancies between the data-generation distributions of synthetic and raw data. Fundamentally, the generational effect underscores a key concern: regardless of the size of synthetic data, generation errors can compromise the accuracy of a statistical method.
@@
While evaluating predictive tasks is typically straightforward, hypothesis testing presents the challenge of regulating the Type-I error. To address this challenge and enhance the power of a test, we introduce “Syn-Test,” a test that augments the sample size of raw data by applying synthetic data. For clarity, “Syn-A” denotes Method A within the Syn framework throughout this article. Syn-Test determines the ideal size of synthetic data required to manage the empirical Type-I error while performing a test for finite inference samples using Monte Carlo methods. Our research indicates that the ideal size of synthetic data can heighten the accuracy. Moreover, our theoretical investigation sheds light on the generational effect, precision, and the size of synthetic data.
Additionally, we introduce a streamlined approach called Syn-Slm to improve Syn’s usability in applications. This approach forgoes actual data generation when one knows the synthetic data distribution. Using sentiment analysis as an illustration, we demonstrate that Syn-Slm is competitive with some alternatives under the Syn framework.
To showcase the capabilities of the Syn framework, we delve into three key domains: sentiment analysis of texts, predictive modeling for structured data, and inference on tabular data. Across these domains, all statistical methods leveraging high-fidelity synthetic data surpass their counterparts employing raw data. The superior performance is due to high-fidelity synthetic data generated by diffusion models. Initially trained on raw data, these models are further improved by fine-tuning pre-trained models via knowledge transfer, resulting in enhanced statistical accuracy and larger sample sizes.
@@
In the first area, we contrasted three models: OpenAI’s Generative Pre-trained Transformer (GPT)-3.5 within the Syn framework, Distilling BERT (DistilBERT, Sanh et al. (2019)) in the Syn-Slm framework, and the Long Short-term Memory Networks (LSTM) in the traditional framework for analyzing consumer reviews from the IMDB movie dataset. In this context, Syn’s generative capability using GPT-3.5 significantly outperforms the LSTM approach. Yet, the Syn-Slm framework using DistilBERT, although trailing, demonstrated notable competitiveness compared to GPT-3.5.
@@
In the second area, we introduced Syn-Boost, a version of CatBoost (Dorogush, Ershov and Gulin, 2018)—a gradient boosting algorithm (Schapire, 1990)—trained on synthetic data. Syn-Boost bolsters the precision of CatBoost for both regression and classification tasks across eight real-world datasets, utilizing a refined tabular diffusion model (Kotelnikov et al., 2023). Statistically, the error trajectory of Syn-Boost exhibits either a U-shaped or L-shaped pattern, determined by the generation errors and the volume of synthetic data used. Moreover, when employing the same knowledge transfer methods, Syn-Boost surpasses traditional feed-forward networks trained on raw data in six of eight cases. These observations highlight that the generative approach of Syn offers a predictive advantage even against a top predictive methodology given the same data input.
@@
In the third domain, we explore the pivotal role of significance tests in discerning feature relevance in regression and classification using CatBoost. This exploration within black-box models unveils largely uncharted territory. Recently, Dai, Shen and Pan (2022) introduced a nonparametric asymptotic test through sample splitting. To augment its statistical prowess, we employ the Syn-Test, capitalizing on pre-trained generative models and knowledge transfer, as demonstrated in two distinct scenarios. In one scenario, we leverage a pre-trained model to ensure smooth knowledge transfer from the male data to the female data while improving generation fidelity and test accuracy for female data, especially when male and female data distributions present distinct characteristics. These observations emphasize the importance of knowledge transfer in mitigating disparities in domains such as healthcare and social science, particularly when data for specific subgroups, like minorities, are limited. Moreover, we shed light on the “generational effect”, accentuating the invaluable interplay of synthetic data generation and knowledge transfer in hypothesis testing—a domain rarely explored in current literature.
@@
This article consists of six sections. Section 2 explores Syn’s role in enhancing the accuracy of statistical methods and emphasizes the importance of knowledge transfer in synthetic data generation. Section 3 discusses the streamlined approach—Syn-Slm and its impact on generative techniques in statistical practice. Section 4 provides illustrative examples, exploring the generational effect in predictive modeling and inference. Section 5 focuses on the privacy issue of synthetic data. In Section 6, we discuss the implications of generating synthetic data for data science. The Appendix contains technical details.
@@ 2.1 Synthetic Data
It is imperative to underscore the pivotal role of knowledge encapsulated in pre-trained generative models for improving generation accuracy through fine-tuning. Nevertheless, directly accessing the pre-training data for pre-trained models is frequently infeasible, impeded by privacy concerns, extensive storage needs, and data inconsistencies (Sweeney, 2002; Quionero-Candela et al., 2009; Labrinidis and Jagadish, 2012; Carroll, 2006), as seen in models like GPT-4. Furthermore, the distributions of pre-training datasets, especially from different sources, might not always mirror that of raw data, as illustrated by the Adult-Male and Adult-Female examples in Section 4.3. Considering these challenges and the nuances of real-world scenarios, we omit pre-training data from our raw data composition throughout this article.
To yield $\tilde{F}$ directly, one can utilize some revertible generative models such as normalizing flows and Roundtrip GAN (Liu et al., 2020), acting as a nonparametric estimate of $F$. For other generative models, such as diffusion models and GPT, one can typically obtain $\tilde{F}$ from synthetic data employing Monte Carlo methods, which we elaborate on in Section 3. In the numerical examples presented in this paper, we utilize a tabular diffusion model TDM (Kotelnikov et al., 2023) and GPT for synthetic data generation.
@@ 2.2 Optimal Synthetic Size for Estimation and Prediction
As stated in Theorem 2.1, a sizable generation error leads to the synthetic risk $\text{R}(\widehat{\bm{\theta}}(\tilde{\bm{Z}}^{(m)}))$ minimizing at a finite $m_{0}$, after which the synthetic error starts to deteriorate. This result suggests that augmented synthetic sample size may not enhance estimation or prediction given substantial generation error, a phenomenon further exemplified in Section 4.2. Conversely, with a small generation error, the synthetic risk remains controlled, and the optimal $m_{0}$ tends to be significantly large or infinite, as substantiated by Theorem 2.2.
@@ 2.3 Optimal Synthetic Size for Hypothesis Testing
Given a raw sample, Syn-Test employs two nearly equal-sized subsamples $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, partitioned from a training sample for fine-tuning a pre-trained generative model. It also uses a separate inference sample $\mathcal{S}_{3}$ of size $n$ for validating model training. One generative model generates synthetic data using $\mathcal{S}_{1}$ for null distribution estimation, while the other uses $\mathcal{S}_{2}$ for computing the test statistic. Figure 4 illustrates the splitting scheme and Syn-Test process. Syn-Test also empirically determines the optimal synthetic data size $m_{0}$ to control the Type-I error. By swapping the roles of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, Syn-Test can de-randomize the partition, transitioning the original inference sample size $n$ to a synthetic inference sample size of $m$. This swapping mechanism proves especially advantageous in scenarios with low generation error. Crucially, abundant synthetic data can enhance the size of the raw data, even when sample splitting results in a reduced inference sample size (Wasserman and Roeder, 2009; Wasserman, Ramdas and Balakrishnan, 2020).
@@
Step 1: Controlling Type-I Error. Generate $D$ distinct synthetic samples $(T(\tilde{\bm{Z}}_{1}^{(m,d)}))_{d=1}^{D}$ of size $m$ by refining a pre-trained generative model with $\mathcal{S}_{1}$ under $H_{0}$. Compute the empirical distribution of the test statistic $T$ using $(T(\tilde{\bm{Z}}_{1}^{(m,d)}))_{d=1}^{D}$. Define a rejection region $C_{m}$ at a significance level where $\alpha>0$.
@@
Step 3: Calculating the P-value. With the determined $\hat{m}$, produce synthetic data $\tilde{\bm{Z}}_{1}^{(\hat{m})}$ by fine-tuning the pre-trained generative model using $\mathcal{S}_{1}$. Calculate the test statistic for $T(\tilde{\bm{Z}}_{1}^{(\hat{m})})$ and determine the P-value, $P^{1}$, leveraging the null CDF based on $\mathcal{S}_{2}$.
@@
Syn-Test enables valid inference without requiring many model assumptions, specific data distributions, and an infinitely large inference sample. Instead, its validity and power hinge on the generation error, which is generally satisfied when using generative models trained on adequately large datasets.
Syn-Test is adept at debiasing test statistics. In this context, the bias estimation is achieved through a Monte Carlo (MC) approach, utilizing synthetic samples generated by a refined generative model. This estimated bias is then subtracted from the estimated null distribution, leading to a debiased version of the test statistic. The effectiveness of this aspect of Syn-Test is demonstrated through numerical examples, as detailed in Section 4.3.

### 10 Mar 2024 v3

@@ Abstract
Synthetic data generation, a cornerstone of Generative Artificial Intelligence, promotes a paradigm shift in data science by addressing data scarcity and privacy while enabling unprecedented performance. As synthetic data becomes more prevalent, concerns emerge regarding the accuracy of statistical methods when applied to synthetic data in contrast to raw data. This article explores the effectiveness of statistical methods on synthetic data and the privacy risks of synthetic data. Regarding effectiveness, we present the Synthetic Data Generation for Analytics framework. This framework applies statistical approaches to high-quality synthetic data produced by generative models like tabular diffusion models, which, initially trained on raw data, benefit from insights from pertinent studies through transfer learning. A key finding within this framework is the generational effect, which reveals that the error rate of statistical methods on synthetic data decreases with the addition of more synthetic data but may eventually rise or stabilize. This phenomenon, stemming from the challenge of accurately mirroring raw data distributions, highlights a “reflection point”—an ideal volume of synthetic data defined by specific error metrics. Through three case studies, sentiment analysis, predictive modeling of structured data, and inference in tabular data, we validate the superior performance of this framework compared to conventional approaches. On privacy, synthetic data imposes lower risks while supporting the differential privacy standard. These studies underscore synthetic data’s untapped potential in redefining data science’s landscape.
@@ 1 Introduction
The advent of synthetic data generation, fueled by generative artificial intelligence (GAI), has initiated a paradigm shift in data analytics, pivoting towards the adoption of synthetic data. Synthetic data, designed to simulate real-world scenarios, presents a viable alternative to the challenges associated with traditional data collection, sharing, and analysis, especially when data are scarce or sensitive. As predicted by Gartner, synthetic data will account for 60% of the data used in AI and analytics projects by 2024, surpassing the utilization of raw data in AI models by 2030 (Gartner, 2022; Eastwood, 2023). Highlighting its transformative impact, a Forbes article by Toews (2022) predicts that synthetic data technology ”will reshape the world of AI in the years ahead, scrambling competitive landscapes and redefining technology stacks.” Furthermore, Jordon et al. (2022) provides an exhaustive review of synthetic data, highlighting its considerable promise as a technology. This evolution signifies a shift away from the traditional reliance on raw data, opening a new chapter in data generation and its application of analytics.
This paper investigates the challenges of efficacy and privacy in synthetic data utilization, emphasizing the role of synthetic data in enhancing data analytics. Specifically, we introduce the Synthetic Data Generation for Analytics (Syn) framework, aimed at increasing the precision of statistical methods applied to high-fidelity synthetic data that accurately replicates raw data through transfer learning. This framework mirrors the statistical properties of raw data through advanced knowledge transfer techniques while offering a potential solution for data sharing without compromising data privacy. Synthetic data provides two main benefits for data analytics: it mitigates data scarcity and addresses privacy issues (Ghalebikesabi et al., 2023).
@@
In the Syn framework, a generative model, informed by raw data, is used to produce synthetic data that replicates its distribution, incorporating insights from pre-trained models in relevant studies through transfer learning. The Syn framework enables the integration of both pre-trained and generative models of various kinds that share comparable latent structures (Akbar, Wang and Eklund, 2023). It employs an array of generative models tailored for different domains: image diffusion (Sohl-Dickstein et al., 2015; Wei et al., 2021), text diffusion models (Ho, Jain and Abbeel, 2020), text-to-image diffusion models (Zhang et al., 2023a), time-series diffusion models (Lin et al., 2023), spatio-temporal diffusion models (Yuan et al., 2023), and tabular diffusion models (Kotelnikov et al., 2023; Zheng and Charoenphakdee, 2022; Kim, Lee and Park, 2022; Lee, Kim and Park, 2023). Moreover, Syn embraces advanced models such as the Reversible Generative Models (Kingma and Dhariwal, 2018). These flow-based models capture the raw data distribution and can estimate both conditional and marginal distributions.
Through Syn exploration, we address a critical issue of efficacy: Can high-fidelity synthetic data boost the effectiveness of statistical methods solely reliant on raw data? If so, how may we implement such enhancements? Recent research offers diverging viewpoints on this issue. In some cases, synthetic X-ray images improve the accuracy of machine learning models (Gao et al., 2023), whereas, in others, training on synthetic data may compromise performance for some machine learning models (Kotelnikov et al., 2023).
@@
Our investigation reveals that synthetic data, when accurately generated, can boost the accuracy of a statistical method by augmenting the sample size of raw data. However, a significant caveat emerges: a statistical method applied to low-fidelity synthetic data could yield unreliable outcomes. Often, a generational effect emerges, whereby as the size of synthetic data grows, the precision gain of a method may diminish or even plateau. This phenomenon has been illustrated in our study for structured data prediction in Section 3.2. This challenge arises from generation errors or discrepancies between the data-generation distributions of synthetic and raw data. Fundamentally, the generational effect underscores a key concern: regardless of the size of synthetic data, generation errors can compromise the accuracy of a statistical method.
@@
While evaluating a supervised task is straightforward, hypothesis testing presents the challenge of regulating the Type-I error while enhancing the power of a test. To address this challenge, we introduce “Syn-Test,” a test that augments the sample size of raw data by applying synthetic data. For clarity, “Syn-A” denotes Method A within the Syn framework throughout this article. Syn-Test determines the ideal size of synthetic data required to manage the empirical Type-I error while performing a test for finite inference samples using Monte Carlo methods. Our research indicates that the ideal size of synthetic data can heighten the accuracy. Moreover, our theoretical investigation sheds light on the generational effect, precision, and the size of synthetic data.
Additionally, we introduce Syn-Slm, a streamlined approach enhancing Syn’s usability in various applications. This method bypasses the training of conventional predictive models in supervised learning tasks. Instead, it directly models the outcome’s conditional distribution through advanced generative models. The effectiveness of Syn-Slm is demonstrated with two examples: sentiment analysis using text data and regression analysis with tabular data.
To highlight the capabilities of the Syn framework, we explore sentiment analysis, predictive modeling for structured data, and tabular data inference. In these domains, methods using high-fidelity synthetic data outperform those with raw data, attributed to the synthetic data generation by diffusion models obtained by fine-tuning relevant pre-trained models on raw data. This fine-tuning boosts efficiency and accuracy by leveraging existing knowledge akin to a scientist solving specialized problems. Specifically, it updates a generative model’s architecture, like adjusting a neural network’s weights or training the entire network or a subnetwork, with the rest remaining unchanged. Furthermore, one may augment network architectures with adapters to accommodate various input dimensions and tasks. For example, a tabular diffusion model pre-trained on an Adult-Male dataset can be fine-tuned on an Adult-Female dataset with the same features, as discussed in Section 3.2.1. Similarly, a diffusion model initially trained with X-ray images (Akbar, Wang and Eklund, 2023) could undergo fine-tuning, including potential modifications to the network architecture, for CT-scan images. Despite originating from distinct domains, they exhibit visual resemblances.
@@
In the first area, we contrast three models: OpenAI’s Generative Pre-trained Transformer (GPT)-3.5 for illustrating Syn-Slm, DistilBERT (Sanh et al., 2019)) using transfer learning, and LSTM in the traditional framework for analyzing consumer reviews from the IMDB movie dataset. In this context, Syn’s generative capability using GPT-3.5 significantly outperforms the LSTM approach in the traditional setup. Yet, DistilBERT in the transfer learning setup, although trailing, demonstrates notable competitiveness compared to GPT-3.5. To further validate Syn-Slm’s application, we provide a simulated example with tabular data in Appendix B.
@@
In the second area, we introduce Syn-Boost, a version of CatBoost (Dorogush, Ershov and Gulin, 2018)—a gradient boosting algorithm (Schapire, 1990)—trained on synthetic data. Syn-Boost bolsters the precision of CatBoost for both regression and classification tasks across eight real-world datasets, utilizing a refined tabular diffusion model (Kotelnikov et al., 2023). Statistically, the error trajectory of Syn-Boost exhibits either a U-shaped or L-shaped pattern, determined by the generation errors and the volume of synthetic data used. Moreover, when employing the same knowledge transfer methods, Syn-Boost surpasses traditional feed-forward networks trained on raw data in six of eight cases. These observations highlight that the generative approach of Syn offers a predictive advantage even against a top predictive methodology given the same data input.
@@
In the third domain, we explore feature significance tests in discerning feature relevance in regression and classification using CatBoost. This exploration within black-box models unveils largely uncharted territory. Recently, Dai, Shen and Pan (2024) introduces an asymptotic test through sample splitting. To augment its statistical prowess, we employ the Syn-Test, capitalizing on pre-trained generative models and knowledge transfer, as demonstrated in two distinct scenarios. In one scenario, we leverage a pre-trained model to ensure smooth knowledge transfer from the male data to the female data while improving generation fidelity and test accuracy for female data, especially when male and female data distributions present distinct characteristics. These observations emphasize the importance of knowledge transfer in mitigating disparities in domains such as healthcare and social science, particularly when data for specific subgroups, like minorities, are limited. Moreover, we shed light on the “generational effect”, accentuating the invaluable interplay of synthetic data generation and knowledge transfer in hypothesis testing—a domain rarely explored in current literature.
@@
This article consists of six sections. Section 2 explores Syn’s role in enhancing the accuracy of statistical methods and emphasizes the importance of knowledge transfer in synthetic data generation. Section 3 provides illustrative examples, exploring the generational effect in predictive modeling and inference. Section 4 focuses on the privacy issue of synthetic data. In Section 5, we discuss the implications of generating synthetic data for data science. The Appendix contains technical details.
@@ 2.1 Synthetic Data
It is imperative to underscore the pivotal role of knowledge encapsulated in pre-trained models for improving generation accuracy through fine-tuning. Nevertheless, directly accessing the pre-training data for pre-trained models is often infeasible, impeded by privacy concerns, extensive storage needs, and data inconsistencies (Sweeney, 2002; Quionero-Candela et al., 2009; Labrinidis and Jagadish, 2012; Carroll, 2006), as seen in models like GPT-4. Furthermore, the distributions of pre-training datasets, especially from different sources, might not always mirror that of raw data, as illustrated by the Adult-Male and Adult-Female examples in Section 3.3. Considering these challenges and the nuances of real-world scenarios, we omit pre-training data from our raw data composition throughout this article.
To yield $\tilde{F}$ directly, one can utilize some revertible generative models such as normalizing flows and Roundtrip GAN (Liu et al., 2020), acting as a nonparametric estimate of $F$. For other generative models, such as diffusion models and GPT, one can typically obtain $\tilde{F}$ from synthetic data employing Monte Carlo methods, which we elaborate on in Section 2.4. In the numerical examples presented in this paper, we utilize a tabular diffusion model (TDM, Kotelnikov et al. (2023)) and GPT for synthetic data generation.
@@ 2.2 Optimal Synthetic Size for Estimation and Prediction
  1. (i)
    The optimal synthetic size $m_{0}<m^{*}$ is finite.
  2. (ii)
[…]
The premise regarding the growth of $Gr^{(m)}$ in $m$ in Theorem 2.1 suggests that the divergence between the synthetic risk and the risk escalates more rapidly when $m\leq m^{*}$, beyond which it remains higher than $Gr^{(m^{*})}$. This condition implies a substantial generation error. According to Theorem 1, a considerable generation error results in the synthetic risk $\text{R}(\widehat{\bm{\theta}}(\tilde{\bm{Z}}^{(m)}))$ reaching its minimum at a specific $m_{0}$, beyond which the synthetic error begins to worsen. This outcome indicates that an increase in synthetic sample size does not necessarily improve estimation or prediction accuracy in the presence of significant generation error, a phenomenon further illustrated in Section 3.2. In contrast, when the generation error is minimal, the synthetic risk remains controlled, and the optimal $m_{0}$ tends to be substantially large or potentially infinite, as suggested by Theorem 2.2.
@@ 2.3 Optimal Synthetic Size for Hypothesis Testing
Given a raw sample, Syn-Test employs two nearly equal-sized subsamples $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, partitioned from a training sample for fine-tuning a pre-trained generative model. It also uses a separate validation sample $\mathcal{S}_{3}$ for validating model training. One generative model generates synthetic data using $\mathcal{S}_{1}$ for null distribution estimation, while the other uses $\mathcal{S}_{2}$ for computing the test statistic. Figure 7 illustrates the splitting scheme and Syn-Test process. Syn-Test also empirically determines the optimal synthetic data size $m_{0}$ to control the Type-I error. By swapping the roles of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, Syn-Test can de-randomize the partition, transitioning the original inference sample size $n$ to a synthetic inference sample size of $m$. This swapping mechanism proves especially advantageous in scenarios with low generation error. Crucially, abundant synthetic data can enhance the size of the raw data, even when sample splitting results in a reduced inference sample size (Wasserman and Roeder, 2009; Wasserman, Ramdas and Balakrishnan, 2020).
@@
Step 1: Controlling Type-I Error. Generate $D$ distinct synthetic samples $(T(\tilde{\bm{Z}}_{1}^{(m,d)}))_{d=1}^{D}$ of size $m$ by refining a pre-trained generative model with $\mathcal{S}_{1}$ under $H_{0}$, using $\mathcal{S}_{3}$ as a validation set for model tuning to avoid overfitting. Compute the empirical distribution of the test statistic $T$ using $(T(\tilde{\bm{Z}}_{1}^{(m,d)}))_{d=1}^{D}$. Define a rejection region $C_{m}$ at a significance level where $\alpha>0$.
@@
Step 3: Calculating the P-value. With the determined $\hat{m}$, produce synthetic data $\tilde{\bm{Z}}_{2}^{(\hat{m})}$ by fine-tuning the pre-trained generative model using $\mathcal{S}_{2}$ with $\mathcal{S}_{3}$ as a validation set. Calculate the test statistic for $T(\tilde{\bm{Z}}_{2}^{(\hat{m})})$ and determine the P-value, $P^{1}$, leveraging the null CDF estimated from $\mathcal{S}_{1}$ in Step 1.
@@
Syn-Test facilitates valid inference without necessitating many model assumptions, specific data distributions, or an infinitely large inference sample. Its validity and power primarily hinge on the small generation error, a condition usually met by generative models trained on sufficiently large datasets. Additionally, Syn-Test permits a fine-tuned generator to estimate the bias of a test statistic through a Monte Carlo approach. This estimation can serve various purposes in downstream analysis, including debiasing. We defer this methodology in the future study.
2.4 Syn-Slm: Streamlined Approach
The Syn Framework enables synthetic data generation, mirroring raw data distributions. It is adept at tackling statistical challenges in both unsupervised and supervised realms. It can also derive $\tilde{F}$ directly from some generative models such as normalizing flows (Dinh, Krueger and Bengio, 2014; Dinh, Sohl-Dickstein and Bengio, 2016; Kingma and Dhariwal, 2018). Subsequently, we introduce Syn-Slm, a streamlined approach for supervised learning tasks. Unlike the methods discussed in Section 2.2, Syn-Slm directly models the conditional distribution of the outcome given predictors without relying on an additional predictive model for synthetic data. This method simplifies the process by directly modeling the conditional distribution of the outcome.
Supervised learning aims to predict an outcome based on a set of predictors, denoted as $\bm{X}$. Consider a scenario with a one-dimensional outcome variable $Y$ and define $\bm{Z}=(Y,\bm{X})$. In this context, the goal is to estimate a statistical functional, represented as $\phi(F_{Y|\bm{X}})$, where $F_{Y|\bm{X}}$ is the conditional CDF of $Y$ given $\bm{X}$. Our Syn-Slm method introduces a plug-in estimate $\phi(\tilde{F}_{Y|\bm{X}})$, where $\tilde{F}_{Y|\bm{X}}$ represents the conditional CDF of a synthetic $Y$ given $\bm{X}$. This approach, focusing on estimating $\tilde{F}_{Y|\bm{X}}$, is notably distinct from conventional methods, which often estimate $\tilde{F}_{\bm{X}|Y}$, such as generating images from a specific class. However, as hypothesized by Zhang et al. (2023b), this type of generative modeling could outperform direct predictive modeling, as demonstrated in their work using imputation techniques. To illustrate, if $\phi(F_{Y|\bm{X}})=\operatorname{E}(Y|\bm{X})$ signifies the conditional expectation of $Y$, then $\phi(\tilde{F}_{Y|\bm{X}})=\operatorname{E}(\tilde{Y}|\bm{X})$, where $\tilde{Y}$ is derived from $\tilde{F}_{Y|\bm{X}}$, resulting in a Syn-Slm estimate. An example is provided in Section 3.1, and a simulated example is presented in Appendix B.
In contrast, the Syn prediction method described in Section 2.2 employs a two-step process: it initially generates joint synthetic data $\tilde{\bm{Z}}=(\tilde{Y},\tilde{\bm{X}})$, followed by training a predictive model using $\tilde{\bm{Z}}$. For instance, Syn-Boost in 3.2 utilizes TDM (Kotelnikov et al., 2023) as the generator and CatBoost (Dorogush, Ershov and Gulin, 2018) for prediction. While straightforward in implementation, methods akin to Syn-Boost may incur prediction errors from predictive modeling and generation errors from generating $\bm{Z}$. In contrast, Syn-Slm directly models the conditional generation of $Y|\bm{X}$ and estimates $\tilde{F}_{Y|\bm{X}}$ via a Monte Carlo approach, offering a fully non-parametric solution. Additionally, Syn-Slm enables the estimation of various statistical functionals $\phi(F_{Y|\bm{X}})$, such as $E(Y|\bm{X})$, $\text{Var}(Y|\bm{X})$, and quantiles of $Y|\bm{X}$. In contrast, methods like Syn-Boost require training different predictive models under varying assumptions to estimate these quantities. A detailed comparison between Syn-Boost and Syn-Slm is presented in Table 1.
Table 1: Comparison between Syn-Boost and Syn-Slm in supervised learning.
Syn-Boost Syn-Slm
Generative modeling Joint distribution of $(Y,\bm{X})$ Conditional distribution of $Y|\bm{X}$
Estimation of $\phi(F_{Y|\bm{X}})$ Predictive modeling on $(\tilde{Y},\tilde{X})$ MC computation with $\tilde{Y}|\bm{X}$
Pros
Generative modeling is easy;
Prediction is fast.
Fully non-parametric;
Estimate different $\phi$ once and for all.
Cons
Additional prediction error;
Needs to tune synthetic size.
Generative modeling is challenging;
Prediction relies on MC simulation.